### v2

# suppSupplementary References

Abstract
Abstract: Raw nanopore signal analysis is a common approach in genomics to provide fast and resource-efficient analysis without translating the signals to bases (i.e., without basecalling). However, existing solutions cannot interpret raw signals directly if a reference genome is unknown due to a lack of accurate mechanisms to handle increased noise in pairwise raw signal comparison. Our goal is to enable the direct analysis of raw signals without a reference genome. To this end, we propose Rawsamble, the first mechanism that can identify regions of similarity between all raw signal pairs, known as all-vs-all overlapping, using a hash-based search mechanism.
We show that an off-the-shelf assembler (i.e., miniasm) can use the overlaps found by Rawsamble to construct genomes from scratch (i.e., de novo assembly). Our extensive evaluations across multiple genomes of varying sizes show that Rawsamble provides a significant speedup (on average by 4.48×4.48$\times$4.48 × and up to 23.10×23.10$\times$23.10 ×) and reduces peak memory usage (on average by 4.33×4.$3$ and up to by 22.00×22.$0$) compared to a conventional genome assembly pipeline using the state-of-the-art tools for basecalling (Dorado’s fastest mode) and overlapping (minimap2) on a CPU. We find that $30.94\%$ of overlapping pairs generated by Rawsamble are identical to those generated by minimap2. We further evaluate the benefits and directions that this new overlapping approach can enable in two ways. First, by constructing and evaluating the assembly graphs from the Rawsamble overlaps, we show that we can construct accurate and contiguous assembly segments (unitigs) up to 2.3 million bases in length (almost half the genome length of E. coli). Second, we identify previously unexplored directions that can be enabled by finding overlaps and constructing de novo assemblies as well as the challenges to tackle these future directions.
Availability and Implementation: Rawsamble is available at https://github.com/CMU-SAFARI/RawHash. We also provide the scripts to fully reproduce our results on our GitHub page.
1 Introduction
Nanopore sequencing technology [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] can sequence long nucleic acid molecules, known as reads, of up to a few million bases at high throughput [24, 25, 26]. As a molecule moves through a tiny pore, called a nanopore, ionic current measurements, called raw signals, are generated [5]. Nanopore sequencing provides three unique key benefits.
First, nanopore sequencing enables stopping the sequencing of single reads or the entire sequencing run early, known as adaptive sampling or selective sequencing [27], while raw signals are generated and analyzed during sequencing, called real-time analysis. Adaptive sampling can substantially reduce the sequencing time and cost by avoiding unnecessary sequencing. Second, raw nanopore signals retain more information than nucleotides, such as methylation, and can be useful for many applications [28]. Third, compact nanopore sequencing devices enable on-site portable sequencing and analysis, which can be coupled with real-time analysis [29].
Existing works that analyze raw nanopore signals [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 27, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66] mainly utilize complex deep learning mechanisms [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] to translate these signals into nucleotides, a process called basecalling. Basecalling mechanisms usually 1) are designed to use large chunks of raw signal data for accurate analysis [53], 2) have high computational requirements [60, 26], and 3) oblivious to the similarity information between raw signals that lead to redundant basecalling on these similar regions. Existing basecalling techniques and the lack of understanding of similarity between raw signals can impose limitations to enable 1) accurate real-time analysis [53] and 2) portable sequencing with constrained resources [60] and prevent future work on designing a new class of basecalling techniques that can utilize the similarity information between raw signals.
To fully utilize the unique benefits of nanopore sequencing, it is necessary to analyze raw signals with 1) high accuracy and low latency for adaptive sampling and 2) low resource usage for portability and efficiency. To achieve this, several mechanisms focus on analyzing raw nanopore signals without basecalling [52, 27, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66] by mapping these signals to a reference genome. [^1] We use the raw signal analysis term specifically for these mechanisms in the remainder of the paper. Although raw nanopore signal mapping is widely studied [27, 53, 54, 57, 58, 59, 60, 62, 63, 64, 65, 66], none of these works without a reference genome to identify similarities directly between reads, called all-vs-all overlapping [67, 26, 68].
Although identifying all-vs-all overlapping between raw nanopore signals can be promising to enable new directions for both basecalling mechanisms (as we discuss in Section 4) and raw signal analysis (e.g., de novo assembly construction as we show in Section 3.4), it is challenging to do so for several reasons [5, 69, 70, 71]. First, similarities between a pair of signals must be identified accurately when both raw signals are noisy as compared to identifying similarities between a noisy raw signal and an accurate signal generated from a reference genome [53, 58, 59, 57]. Converting reference genomes to their expected raw signal values is free from certain types of noise that raw nanopore signals contain (e.g., stochastic signal fluctuations [5], variable speed of DNA molecules moving through nanopores [69, 70], and raw reads containing multiple split molecules), while raw signals are not free from such noise.
Second, existing raw signal analysis works lack the mapping strategies typically used in all-vs-all overlapping, such as reporting multiple mappings (i.e., overlaps) of a read to many reads. This is because these works are mainly designed to stop the mapping process as soon as there is an accurate mapping for a read to minimize the unnecessary sequencing [54]. For all-vs-all overlapping, pairwise mappings of a read to many reads (instead of a single mapping) must be reported while avoiding certain trivial cyclic pairwise mappings to construct an assembly [67].
Third, read overlapping can increase the space requirements for storing and using indexing. This is because a read set is likely to be sequenced such that its overall number of bases is larger than the number of bases in its corresponding genome. Such an increased index size can raise the computational and space demands for read overlapping. It is essential to provide high-throughput and scalable computation to enable real-time downstream analysis for future work (as discussed in Section 4).
Our goal is to enable 1) raw signal analysis without a reference genome and 2) new use cases with raw nanopore signal analysis by addressing the challenges of accurate and fast all-vs-all overlapping of raw nanopore signals. To this end, we propose Rawsamble, the first mechanism that enables fast and accurate overlap finding between raw nanopore signals. The key idea in Rawsamble is to re-design the existing state-of-the-art hash-based seeding mechanism [58, 59] for raw signals with more effective noise reduction techniques and useful outputting strategies to find all overlapping pairs accurately, which we explain in three key steps.
First, to enable identifying similarities between a pair of noisy raw signals accurately, Rawsamble filters raw signals to select those substantially distinct from their surrounding signals. Such non-distinct and adjacent signals are usually the result of a certain error type in the analysis, known as stay errors [63, 58, 53]. Although similar filtering strategies [63, 58, 53] are exploited when mapping raw signals to reference genomes, Rawsamble performs a more aggressive filtering to avoid storing the erroneous portions of signals in the index to enable accurate similarity identification from the sufficiently distinct regions of signals. Second, to find multiple overlaps for a read, Rawsamble identifies highly accurate chains from seed matches based on their chaining scores and reports all of these chains as mappings, as opposed to choosing solely the best mapping as determined by weighted decisions among all such chains [59]. Third, to prevent trivial cycles between a pair of overlapping reads, Rawsamble ensures that only one of the overlapping reads in each pair is always chosen as a query sequence, and the other is always chosen as a target sequence based on a deterministic ordering mechanism between these reads. These steps enable Rawsamble to find overlaps between raw signals accurately and quickly.
Rawsamble makes the following key contributions:
  • We propose the first mechanism that can find all-vs-all overlapping of raw signals without basecalling to enable new use cases for raw signal analysis, such as assembly from overlapping raw signals and improving existing basecalling mechanisms.
  • We show that identifying overlaps with Rawsamble 1) is faster (on average 4.48×4.48$\times$4.48 × and up to 23.10×23.10$\times$23.10 ×) and 2) reduces peak memory requirements (on average 4.33×4.$3$ and up to 22.00×22.$0$) compared to the computational pipeline that includes the state-of-the-art basecalling running on a CPU with its fastest model followed by minimap2 to find overlaps. We evaluate the trade-offs between speed, accuracy, and hardware resources by running the basecaller on a CPU and GPU with fast and high-accuracy models.
  • We show that $30.94\%$ of overlapping pairs that Rawsamble generates are identical to the overlapping pairs generated by minimap2.
  • We report the first de novo assembly graphs ever constructed directly from raw signal overlaps without basecalling. We show that we can construct long unitigs up to 2.3 million bases in length for E. coli, which constitutes half the length of the genome. We find that these assemblies constructed from raw signal overlaps are either as accurate or more accurate than those created by minimap2 overlaps.
  • We identify new use cases that can be enabled by using overlapping raw signals and constructing assemblies from raw signals and discuss the advantages of generating them, such as for improving the accuracy and performance of basecalling. We discuss current limitations and challenges to enable these new use cases as future work.
2 Methods
2.1 Overview
Rawsamble is a mechanism to find overlapping pairs between raw nanopore signals (i.e., all-vs-all overlapping), which can be used by existing assemblers to construct de novo assembly graphs without basecalling, as shown in Figure 1. To achieve this, Rawsamble builds on the state-of-the-art raw signal mapper, RawHash2 [58, 59]. We provide the overview of RawHash2 in Supplementary Section A, where we explain the details related to processing raw signals such as generating events after the segmentation and generating hash values from these segmented raw signals (i.e., events). In the remainder of the paper, we use raw signals to refer to these events generated after segmentation.
Rawsamble extends RawHash2 to support all-vs-all raw signal overlapping in four key steps. First, to enable efficient and accurate indexing from the noisy raw signals ( 1), Rawsamble aggressively filters the raw input after the signal-to-event conversion to avoid nanopore-related errors. Second, to improve the accuracy of overlapping ( 2), Rawsamble adjusts the minimum chaining score to avoid false chains, ensuring that only high-confidence overlaps are considered. Third, to enable finding useful and long connections from many overlaps, Rawsamble adjusts the output strategy such that 1) all chains rather than only the best chain are reported and 2) cyclic overlaps are avoided. Fourth, Rawsamble enables the use of existing de novo assemblers off-the-shelf, such as miniasm [67], by providing the overlap information in a standardized format that these assemblers use.
Figure 1: Overview of Rawsamble. We use colors for the inputs, steps, and outputs to highlight the parts that Rawsamble modifies over RawHash2.
2.2 Constructing an Index from Noisy Raw Signals via Aggressive Filtering
Rawsamble identifies overlapping regions between raw nanopore signals using a hash-based seeding mechanism that operates in two steps. First, to enable quick matching between raw signals, Rawsamble enables constructing an index directly from raw nanopore signals instead of using an existing reference genome sequence. While converting reference genomes to their expected raw signal values is mainly free from certain types of noise that raw nanopore signals contain (e.g., stochastic signal fluctuations [5] and variable speed of DNA molecules moving through nanopores [69, 70]), noise in raw nanopore signals can cause challenges to find accurate matches between raw signals when these signals are stored in an index. Second, to reduce noise stored in the hash tables and enable accurate similarity identification when both signals are noisy, Rawsamble aggressively filters raw signals as shown in Figure 2. The filtering mechanism iteratively compares two adjacent signals, $s_{i}$ and $s_{j}$, and removes the second signal, $s_{j}$, if the absolute difference between the two adjacent signals is below a certain threshold $T$. This filtering generates a list of filtered signals that aggressively aims to reduce the impact of stay errors during sequencing. Although similar filtering approaches are used in prior works [53, 58, 59] to reduce the stay errors (e.g., by aiming to perform homopolymer compression of raw signals [63]), Rawsamble employs a substantially larger threshold (i.e., aggressive) for filtering to minimize noise both in the index and during mapping. By applying the aggressive filtering technique, Rawsamble ensures that only high-quality, informative events are used in indexing to improve the accuracy and efficiency of the overall overlapping mechanism when both signals are noisy.
Figure 2: Filtering in Rawsamble. Values in gray boxes show the filtered signals as their values are close to the previous signal (in a blue box) that is not filtered out.
2.3 Adjusting the Chaining Mechanism for Overlapping
To reduce the number of chains that do not result in mapping (i.e., false chains) and construct longer chains, Rawsamble adjusts the chaining mechanism [72, 59] in two ways. First, Rawsamble constructs chains between seed matches (i.e., anchors) with longer gaps by adjusting the maximum gap length between anchors. This adjustment is needed because the filtering mechanism results in a sparser list with potentially long gaps between signals. Second, to ensure that only high-confidence chains are considered as overlaps among many pairwise overlapping regions, Rawsamble sets a higher minimum chaining score for overlapping than mapping, which effectively filters out spurious matches. These adjustments to the chaining mechanism enable Rawsamble to construct long and accurate chains between noisy and sparse raw signals, improving the overall sensitivity of the overlapping process and further downstream analysis such as de novo assembly construction.
2.4 Adjusting the Mapping Strategy
Rawsamble adjusts the mapping strategy used in RawHash2 in two ways. First, Rawsamble generates pairwise mappings of a read to many reads (instead of a single mapping per read). To achieve this, Rawsamble identifies all valid chains based on their chaining scores and reports all such chains between raw signals. This enables the overlapping of a single raw signal with multiple other raw signals. Second, Rawsamble filters out trivial cyclic overlaps between two reads. To do so, Rawsamble avoids reporting overlapping signal pairs both as query and target in the mapping output, which can complicate the de novo assembly construction process [67, 73, 74]. To avoid these trivial cyclic overlaps, Rawsamble implements a pre-defined and deterministic ordering between raw signals (e.g., based on the lexicographic ordering of read names). By comparing raw signals deterministically, Rawsamble guarantees that only one of each pair of overlapping reads is processed as a query sequence while the other is treated as a target sequence. Reporting all chains while avoiding cyclic overlaps between raw signals allows for a comprehensive representation of the overlapping regions, which is useful for constructing accurate and long assemblies.
Output Format. Rawsamble provides the overlapping information between raw signals using the Pairwise mApping Format (PAF) [67]. To construct an assembly graph from the overlapping information that Rawsamble generates, any assembler that takes PAF files as input can be used, including miniasm [67].
3 Results
3.1 Evaluation Methodology
We implement the improvements we propose in Rawsamble on the RawHash2 implementation [59]. Similar to RawHash2 and minimap2 [72], Rawsamble provides the mapping information in the standard Pairwise mApping Format (PAF) [67]. We basecall the signals on two different hardware setups using Dorado’s 1) high accuracy (HAC) model with an NVIDIA RTX A6000 GPU [75], and 2) fast (Fast) and HAC models with an Intel Xeon Gold 6226R CPU [76]. We include Dorado’s results on a CPU to provide a fair comparison on the same hardware platform, as providing the GPU implementation for Rawsamble is a future work. We use POD5 files for all datasets, as suggested by Dorado [35] for optimal performance. Since no prior works can overlap reads using raw signals, we compare Rawsamble to minimap2, the current state-of-the-art read overlapper for basecalled sequences using the datasets shown in Table 1. Although the R9.4.1 nanopore chemistry is deprecated, these raw datasets are commonly stored and basecalled for downstream analysis. We include datasets from both chemistries to show the feasibility and limitations of Rawsamble for each chemistry. For the R10.4.1 dataset, we use only the raw signals that contain a single molecule (i.e., non-chimeric multiple molecules can be contained in a single nanopore raw signal and later split by basecallers into multiple reads).
Table 1: Details of datasets used in our evaluation.
Organism Chemistry Reads Bases Avg. Read Estimated SRA
Model (#) (#) Length Coverage ($\times$) or DOI
D1 E. coli R9.4 353,948 2,332M 6,588 445$\times$ ERR9127551
D2 Yeast R9.4 50,023 385M 7,698 32$\times$ SRR8648503
D3 Green Algae R9.4 30,012 622M TBC 5.6$\times$ ERR3237140
D4 E. coli R10.4.1 6,412,132 9,400M 1,466 1,797$\times$ 10.26188/25521892
Base counts in millions (M). Coverage is estimated using corresponding reference genomes.
All of the datasets are basecalled using Dorado (HAC).
We evaluate computational requirements in terms of overall run time with 64 CPU threads and peak memory usage. We use 64 CPU threads since the average thread utilization of CPU-based basecalling is around 64. We report the elapsed time, CPU time (when using only CPUs without GPUs), and peak memory usage of 1) Rawsamble and 2) basecalling followed by minimap2 (Dorado + Minimap2). For Rawsamble, we additionally report throughput (average number of signals analyzed per second per single CPU thread) to provide insights about its capabilities for real-time analysis. To measure performance and peak memory usage, we use the time -v command in Linux when running Rawsamble, minimap2, and Dorado. For average speedup and memory comparisons of Rawsamble against other methods, we use the geometric mean to reduce the impact of outlier data points on the average calculation.
We evaluate Rawsamble based on two use cases: 1) read overlapping and 2) de novo assembly by constructing assembly graphs. We generate read overlaps from 1) raw signals using Rawsamble and 2) basecalled sequences (with Dorado’s HAC model) of corresponding strands of raw signals using minimap2. To evaluate the accuracy of all-vs-all overlapping, we calculate the ratio of overlapping pairs 1) shared by both Rawsamble and minimap2, 2) unique to Rawsamble, and 3) unique to minimap2. For de novo assembly, we use off-the-shelf miniasm [67] to generate assembly graphs from the read overlaps that Rawsamble and minimap2 generate. We use miniasm because it enables us to evaluate the impact of overlaps that Rawsamble and minimap2 find by using the same assembler for both of them. To identify the roofline in terms of the assembly contiguity given a dataset, we use a highly accurate assembler as gold standard for our evaluations. To do this, we use Dorado’s most accurate model (i.e., SUP) to construct highly accurate reads and use Flye [77] to construct assemblies from these reads, as suggested in the guidelines by ONT [78]. This approach enables us to evaluate the gap between the assemblies that Rawsamble generates and the golden standard assembly that can be constructed from the same datasets.
To evaluate the contiguity of assemblies constructed from the overlaps that Rawsamble and minimap2 generate, we use several metrics. These metrics are 1) the total length of unitigs (Total Length), 2) the number of nucleotides in the largest assembly graph component (Largest Comp.), 3) unitig length at which 50% of the assembly is covered by unitigs of equal or greater length (N50), 4) area under the Nx curve to generate a more robust evaluation of contiguity (auN) [79], 5) longest unitig length, and 6) the number of unitigs. We estimate the basecalled read length of raw signals based on the sample frequency (i.e., the number of signals produced every second) and the translocation speed (i.e., the number of bases estimated to pass through a nanopore in a single second). Estimating the read length based on these parameters is a common and relatively consistent approach widely adopted in previous related works [54, 53, 58, 59]. We use Bandage [80] to visualize these assemblies. We note that our contiguity evaluations for Rawsamble and minimap2 are based on the assembly graphs constructed by miniasm.
[…]
We provide the parameter settings and versions for each tool as well as the details of the preset parameters in Rawsamble in Supplementary Tables S2 (parameters), S3 (details of presets), and S4 (versions). We provide the scripts to fully reproduce our results on the GitHub repository at https://github.com/CMU-SAFARI/RawHash, which contains the corresponding release version of Rawsamble we show in Supplementary Table S4. We provide the detailed reasoning behind the choice of Flye in order to generate the gold standard in Supplementary Material B.
3.2 Performance and Memory
Throughput. Table 2 shows the throughput (i.e., signals processed per second per CPU thread) that Rawsamble reports. Our goal is to estimate the number of CPU threads needed to achieve a throughput faster than the throughput of a single sequencer. Using as few CPU threads as possible is useful to 1) provide better scalability (i.e., analyzing a larger amount of data with the same computation capabilities) and 2) reduce the overall computational requirements and corresponding energy consumption (i.e., analyzing the same amount of data with less computational capabilities). The latter is especially essential for resource-constrained devices (e.g., devices with external and limited batteries). The throughput of a single nanopore is around 5,000 signals per second [35], and the entire sequencer is usually equipped with 512 nanopores. This means a single sequencer’s throughput is around $2{,}560{,}000$ signals per second ($5{,}000\times 512$[81]. We find that Rawsamble provides an average throughput of $1{,}496{,}148$ signals per second per CPU thread. This means Rawsamble can achieve an analysis throughput faster than a single sequencer’s throughput by using an average of two CPU threads. This shows that the overlapping mechanism of Rawsamble is fast enough for performing real-time overlapping tasks using very few CPU threads.
Table 2: Rawsamble Throughput (signals processed per second per CPU thread).
D1 D2 D3 D4
E. coli Yeast Green Algae E. coli (R10.4.1)
Throughput 1,894,434 1,304,910 1,668,065 1,117,183
Computational resources. Table 3 shows the computational resources of the different approaches. Inside the parentheses provided with Dorado + Minimap2 results, we provide the ratio between the reported result and the corresponding Rawsamble result (if higher than 1×1$\times$1 ×, Rawsamble is better).
Table 3: Comparison of various tools across different organisms in terms of elapsed time, CPU time, and peak memory usage. The values in parentheses represent the ratio of the result shown in the cell compared to the corresponding result of Rawsamble (values higher than 1×1$\times$1 × indicate that Rawsamble performs better). We highlight the cells that tools provide a better and worse results with colors.
Organism Tool Elapsed time CPU time Peak
(hh:mm:ss) (sec) Mem. (GB)
D1 Rawsamble 0:52:07 184,938 7.41
E. coli Dorado CPU (Fast) + Minimap2 2:56:45 (3.39×3.39$\times$3.39 ×) 593,764 (3.21×3.21$\times$3.21 ×) 36.73 (4.96×4.96$\times$4.96 ×)
Dorado CPU (HAC) + Minimap2 13:31:19 (15.57×15.57$\times$15.57 ×) 1,408,712 (7.62×7.62$\times$7.62 ×) 62.34 (8.41×8.41$\times$8.41 ×)
Dorado GPU (HAC) + Minimap2 0:26:48 (0.51×0.51$\times$0.51 ×) NA 26.73 (3.61×3.61$\times$3.61 ×)
D2 Rawsamble 0:01:21 3,753 6.68
Yeast Dorado CPU (Fast) + Minimap2 0:31:11 (23.10×23.10$\times$23.10 ×) 98,558 ($\times26$) 146.95 (22.00×22.$0$)
Dorado CPU (HAC) + Minimap2 2:30:28 (111.46×111.46$\times$111.46 ×) 548,495 (146.15×146.15$\times$146.15 ×) 290.85 (43.54×43.54$\times$43.54 ×)
Dorado GPU (HAC) + Minimap2 0:02:00 (1.48×1.48$\times$1.48 ×) NA 5.98 (0.90×0.90$\times$0.90 ×)
D3 Rawsamble 0:06:20 21,769 10.87
Green Algae Dorado CPU (Fast) + Minimap2 0:30:17 (4.78×4.78$\times$4.78 ×) 103,102 (4.74×4.74$\times$4.74 ×) 38.18 (3.51×3.51$\times$3.51 ×)
Dorado CPU (HAC) + Minimap2 7:29:20 (70.95×70.95$\times$70.95 ×) 178,106 (8.18×8.18$\times$8.18 ×) 20.76 (1.91×1.91$\times$1.91 ×)
Dorado GPU (HAC) + Minimap2 0:03:04 (0.48×0.48$\times$0.48 ×) NA 10.08 (0.93×0.93$\times$0.93 ×)
D4 Rawsamble 16:04:39 2,672,608 112.31
E. coli Dorado CPU (Fast) + Minimap2 17:16:52 (1.07×1.07$\times$1.07 ×) 3,478,136 (1.30×1.30$\times$1.30 ×) 103.1 (0.92×0.92$\times$0.92 ×)
(R10.4.1) Dorado CPU (HAC) + Minimap2 135:05:47 (8.40×8.40$\times$8.40 ×) 19,051,635 (7.13×7.13$\times$7.13 ×) 103.1 (0.92×0.92$\times$0.92 ×)
Dorado GPU (HAC) + Minimap2 4:59:49 (0.31×0.31$\times$0.31 ×) NA 103.1 (0.92×0.92$\times$0.92 ×)
We make three key observations. First, Rawsamble provides a substantial speedup and lower peak memory usage compared to Dorado’s fast model on a CPU followed by minimap2 (Dorado CPU (Fast) + Minimap2) by, on average, 4.48×4.48$\times$4.48 × (elapsed time), 4.77×4.$7$ (CPU time), and 4.33×4.$3$ (peak memory usage). When using Dorado’s HAC model on a CPU, Rawsamble provides even better results on average by 31.89×31.89$\times$31.89 × (elapsed time), 15.96×15.96$\times$15.96 × (CPU time), and 5.03×5.03$\times$5.03 × (peak memory usage), since running the HAC model is computationally more costly than running the fast model.
Second, GPU-based basecalling followed by minimap2 takes 0.58×0.58$\times$0.58 × of the time that Rawsamble takes while requiring higher peak memory usage by 1.29×1.29$\times$1.29 × compared to Rawsamble. Although comparing the results between CPUs and GPUs is not ideal due to the massive parallelism that GPUs provide compared to CPUs, Rawsamble still provides comparable performance given the limited parallelism it uses on CPUs (i.e., 64 threads) compared to the massive parallelism that GPUs provide (e.g., a few tens of thousands of threads). This shows that Rawsamble can provide even faster results when accelerated with GPUs based on its comparison when basecalling is done using CPUs and GPUs, which we leave as future work.
We conclude that Rawsamble can perform read overlapping with higher throughput that can be useful when focusing on real-time analysis and better computational resources than CPU-based basecalling followed by minimap2. We find that Rawsamble’s speed is mainly dependent on the amount of bases stored in the hash table, as the speed decreases with increasing the number of locations that need to be analyzed in the index per read. To resolve this scalability issue, future work should focus on designing indexing and filtering methods that provide a limitation on the volume of signals stored in the index and processed during the overlapping step. These results can also be useful when basecalling raw signals using GPUs to reduce the computational overhead that GPU-based basecalling requires, which we discuss in Section 4.
3.3 All-vs-All Overlapping Statistics
Table 4 shows the all-vs-all overlapping statistics between Rawsamble and minimap2. We make two key observations. First, on average, $30.94\%$ of the overlap pairs generated by Rawsamble are shared with the overlap pairs that minimap2 generates. This shows that a large fraction of overlapping pairs does not require basecalling to generate identical overlapping information identified by minimap2 after basecalling. Second, although Rawsamble can find a substantial amount of overlap pairs identical to the overlaps minimap2 finds, there are still overlapping pairs unique to Rawsamble ($9.67\%$ on average) and minimap2 ($59.39\%$ on average). These differences are likely due to differences in 1) certain parameters (e.g., chaining scores) and 2) increased noise inherent in raw signals compared to basecalled sequences, which can become more pronounced in genomes with greater size, complexity, or repetitiveness. Although maximizing the shared overlaps can provide insights regarding the accuracy of overlapping information that Rawsamble finds, it is not necessary to provide near-identical shared overlap statistics for certain use cases such as constructing de novo assemblies [82]. Instead, contiguous and more accurate assemblies can still be constructed using a smaller but useful portion of shared overlapping pairs [82]. We conclude that Rawsamble provides a mechanism that shares a large portion of overlaps with minimap2, while the shared portions decrease due to the differences in parameters and the noise in raw signal datasets.
Table 4: All-vs-all Overlapping Statistics. Percentages show the overlapping pairs that are 1) unique to Rawsamble, 2) unique to minimap2, and 3) reported in both tools (Shared Overlaps).
Organism Unique to Unique to Shared
[…]
D1 E. coli 8.33 50.62 41.05
D2 Yeast 26.54 28.62 44.84
D3 Green Algae 3.76 78.64 17.61
D4 E. coli (R10.4.1) 0.06 79.69 20.25
3.4 Contiguity and Estimated Accuracy of the Assembly Graphs
To evaluate the impact of overlaps that Rawsamble and minimap2 find, we construct de novo assembly graphs from these overlaps by using miniasm off-the-shelf. Table 5 shows the contiguity statistics and estimated accuracy of these de novo assembly graphs (please see Section 3.1 for our explanation and discussion related to using these accuracy ratios only for comparison purposes, not for calculating the actual assembly identities). We note that the assembly graphs constructed using miniasm provide the connection information used to assemble unitigs, as well as the connections between unitigs that make up subgraph components in the assembly graph. Since we use the miniasm assembler without any modifications, we cannot generate the assembled sequences from raw signals as miniasm supports generating assembled sequences from basecalled sequencing reads. This mainly prevents us from evaluating the actual identity of these assemblies from raw signals, which we further discuss in Section 4.
Table 5: Assembly Statistics. Chained Read Percentage provides our estimate in terms of assembly graph accuracy. Other columns provide our evaluations in terms of assembly contiguity. For the accuracy, we highlight the cells that tools provide a better and worse results with colors.
Dataset Tool Chained Read Largest N50 auN Longest Unitig
Percentage (%) Comp. (bp) (bp) (bp) Unitig (bp) Count
D1 Rawsamble 34.0 4,841,669 1,535,079 1,187,229 2,347,310 32
E. coli minimap2 24.3 5,207,206 5,204,754 5,194,495 5,207,206 4
Gold standard NA 5,235,343 5,235,343 5,235,343 5,235,343 1
D2 Rawsamble 43.5 362,050 41,118 48,106 161,883 396
Yeast minimap2 43.2 1,611,876 134,050 137,172 464,054 282
Gold standard NA 11,835,059 640,934 623,210 1,073,346 68
D3 Rawsamble 41.3 448,422 93,111 107,914 252,038 50
Green Algae minimap2 53.2 198,709 63,310 91,360 198,709 55
Gold standard NA 2,255,807 452,774 538,136 1,667,975 420
D5 Rawsamble 34.6 22,921,373 27,460 43,505 410,905 2,194
E. coli minimap2 16.3 27,250,402 24,111 814,807 5,117,413 2888
(R10.4.1) Gold standard NA 5,230,610 5,230,610 5,230,610 5,230,610 1
First, we find that we can construct long unitigs from the raw signal overlaps that Rawsamble finds. The unitigs we can construct from these raw signal overlaps are substantially longer than the average read length of their corresponding datasets (e.g., the longest unitig length in D1 is 413.25×413.25$\times$413.25 × longer than the average read length of the D1 dataset). Rawsamble is the first work that enables de novo assembly construction directly from raw signal overlaps without basecalling, which has several implications and can enable future work, as we discuss in Section 4.
Second, we observe that the unitigs generated from Rawsamble overlaps are usually less contiguous compared to those generated from minimap2 overlaps, based on all the metrics we show in Table 5. Compared to the gold standard assemblies generated using highly accurate basecalled reads and a state-of-the-art assembler, we find that Rawsamble can still achieve a significant portion of the assembly contiguity, especially for the D1 dataset (E.coli). For example, Rawsamble constructs unitigs with the longest unitig length of 2.3 million bases, which is almost half the length of the E. coli genome, whereas the gold standard assembly produces a single unitig covering the entire genome. This indicates that Rawsamble can achieve substantial contiguity relative to the gold standard without basecalling.
Third, in terms of the accuracy estimate of the assembly graphs (i.e., Chained Read Percentage), we find that a larger percentage of unitig reads generated by the Rawsamble overlaps appear in close proximity when these reads are mapped to their corresponding reference genome compared to the unitig reads generated by the minimap2 overlaps. Although this observation by itself does not show that the Rawsamble overlaps can lead to assemblies with higher identity, it provides the accuracy estimate that these assemblies share similar accuracy to those generated by the minimap2 overlaps.
We conclude that Rawsamble generates useful overlapping information, enabling the construction of assemblies directly from raw signals. The assembly graphs generated by the Rawsamble overlaps mainly provide comparable contiguity and accuracy compared to those generated by minimap2. We discuss the potential next steps to enable evaluating the accuracy of these assemblies in Section 4.
4 Discussion and Future Work
Limitations. Our evaluation demonstrates that Rawsamble achieves high throughput when finding overlaps between raw signal pairs, making it a viable candidate for real-time analysis. However, there are still two main challenges to fully utilize real-time de novo assembly construction during sequencing. First, the hash-based index should be constructed and updated dynamically in real-time while sequencing is in progress. Although Rawsamble provides the mechanisms for storing multiple hash tables that can be constructed for each chunk of raw signals generated in real-time, it is not computationally feasible to dynamically update the index for all sequenced signals as the memory and computational burden increase with each hash table generated. Such an approach requires a decision-making mechanism to dynamically stop updating the index after sequencing a certain amount of signals. The stopping mechanism should be accurate enough to ensure that the current state of the index provides sufficient information to find the overlaps between the already sequenced signals and the new signals generated after the index construction. Second, the assembly graph should be constructed and updated dynamically while the new overlap information is generated in real time. This is challenging as the intermediate steps for generating the unitigs (e.g., the transitive reduction step [83]) in assembly graphs may not work optimally without the full overlap information, as it is likely to remove graph connections that can be useful with new overlap information. It might be feasible to use graph structures that are more suitable for streaming data generation, such as de Bruijn graphs, for assembly construction purposes [84, 85, 86].
Rawsamble can find overlaps only between reads coming from the same strand, as it lacks the capability to construct the reverse complemented version of the signals to identify the matches on the other strands. This potentially leads to gaps in the assembly and construction of the unitigs from both strands. Designing a mechanism that can reverse complement raw signals without basecalling is future work.
Challenges for constructing and evaluating de novo assemblies. Although the main focus of our work is enabling all-vs-all overlapping between raw signals, we identify de novo assembly construction as a natural use case that can utilize the all-vs-all overlap information from raw signals that Rawsamble finds. To show that the overlaps from Rawsamble can lead to unitigs that are much longer than average read lengths, we construct and evaluate the de novo assembly graphs from the Rawsamble overlaps, as discussed in Section 3.4. However, we identify two key challenges that limit the scope of our evaluation of these de novo assemblies, which we leave as future work.
First, we cannot generate the sequence of assembled signals from the assembly graphs (i.e., spelling the sequence from the assembly graph). This is because we use the miniasm assembler to construct these assembly graphs from the raw signal overlaps. Although miniasm can use the overlap position information to construct the assembly graph, it cannot output the sequence of assemblies, as it is mainly designed for basecalled sequences. This is challenging because each raw signal contains metadata used in basecalling mechanisms (e.g., nanopore-specific information such as the baseline current levels). Carrying the metadata information effectively from many overlapping signals is essential to enable further analysis of these assembled signals.
Second, to evaluate the identity of the constructed assemblies, these assemblies need to be aligned to a ground truth assembly. This can be done either by either 1) basecalling the assembled signal or 2) aligning the assembled signal using dynamic time warping (DTW) [57] to the ground truth assembly. These approaches require constructing the assembled signal (i.e., the first challenge mentioned above) and designing a new class of basecallers to directly basecall assembled signals, which we discuss as future work below. Although the missing identity evaluations of these assemblies can be a major limitation of our work, we believe the assemblies that are constructed from the Rawsamble overlaps are likely to be close to the assembly accuracy that can be achieved by using the minimap2 overlaps based on our estimated accuracy evaluation.
New Directions for Future Work. We identify several new directions for future work. First, integrating the overlapping information with basecalling can provide accuracy benefits for basecallers. Existing basecallers are designed to basecall individual reads without such overlapping information. Such a combined approach can provide additional useful features for the underlying machine learning models in basecallers to improve the basecalling accuracy. Second, de novo assembly construction enables identifying the reads that do not provide useful information for assembly if they are fully contained by other overlapping read pairs [67]. Identifying these potentially useless reads early on provides the opportunity to avoid basecalling them, which can improve the overall execution time of basecalling by filtering out such reads. Third, de novo assembly construction from raw signals provides new opportunities to design new basecallers that can basecall the assembled signals. Such a basecaller has the potential to substantially improve the performance, as it enables basecalling fewer long unitigs instead of many shorter reads.
While Rawsamble enables new directions in raw signal analysis, particularly for de novo assembly, several challenges and opportunities for future work remain. Addressing these challenges has the potential to enable new use cases and applications in raw signal and basecalled sequence analyses.
5 Conclusion
We introduce Rawsamble, the first mechanism that can find overlaps between two sets of raw nanopore signals without translating them to bases. We find that Rawsamble can 1) find overlaps while meeting the real-time requirements with an average throughput of $1{,}496{,}148$ signals/sec per CPU thread, 2) reduce the overall time needed for finding overlaps (on average by 4.48×4.48$\times$4.48 × and up to by 23.10×23.10$\times$23.10 ×) and peak memory usage (on average by 4.33×4.$3$ and up to by 22.00×22.$0$) compared to the time and memory needed to run state-of-the-art read mapper (minimap2) combined with the Dorado basecaller running on a CPU with its fastest model, 3) share a large portion of overlapping pairs with minimap2 ($30.94\%$ on average), and 4) construct long assemblies from these useful overlaps. We find that we can construct assemblies of half the length of the entire E. coli genome (i.e., of length 2.3 million bases) directly from raw signal overlaps without basecalling. Finding overlapping pairs from raw signals is critical for enabling new directions that have not been explored before for raw signal analysis, such as de novo assembly construction from overlaps that we explore in this work. We discuss many other new directions that can be enabled by finding overlaps and constructing de novo assemblies.
We hope and believe that Rawsamble enables future work in at least two key directions. First, we aim to fully perform end-to-end genome analysis without basecalling. This can be achieved by further improving tools such as Rawsamble to construct de novo assemblies directly from raw signals such that these assemblies are consistently better than those generated from basecalled sequences in all cases. Second, we should rethink how we train and use modern neural network-based basecallers by integrating additional useful information (e.g., overlaps or assemblies of signals) generated by Rawsamble into these basecallers.
Supplementary Material for
Rawsamble: Overlapping Raw Nanopore Signals
using a Hash-based Seeding Mechanism
A RawHash2 Overview
[…]
Second, to further mitigate this noise issue in events, RawHash2 quantizes the event values such that slightly different event values can be quantized into the same value (i.e., bucketing) to enable direct matching of quantized event values between a reference genome and raw signals. To enable an accurate quantization, RawHash2 identifies the range of values that are assigned to the same quantized value dynamically according to the nanopore model.
Third, to reduce the number of potential matches without reducing accuracy, RawHash2 concatenates the quantized event values of consecutive events (i.e., consecutive k-mers) and generates a hash value from these concatenated values.
[…]
Figure S1: Overview of RawHash2.
B Generating a Gold Standard Assembly
To generate a gold standard assembly using R9.4 Simplex reads, we explore two approaches.
[…]
For the high-coverage D1 E.coli dataset, Hifiasm outputs a fragmented assembly, while Verkko requires extensive parameter tuning (specifically in terms of coverage, ---unitig-abundance, and ---base-k) to achieve desirable contiguity and completeness. LJA produces an almost perfect assembly when a minimum length filter of 30kbp is applied to the corrected reads, as suggested by the HERRO paper. Flye performs comparably to LJA without the need for error correction. It should be noted that the golden standard assembly for D4 E.coli dataset was limited in terms of assembly coverage, using ---asm-coverage 50, as Flye raised an error regarding an otherwise too high coverage level.
For the other datasets, with or without error correction, most assemblers either fail to produce assemblies or generate suboptimal results, except for Flye. We attribute this to their sensitivity to inaccuracies in the uncorrected reads and the reduced coverage after correction. Supplementary Table S1 shows how the coverage levels change after correction for each dataset. Notably, Flye works well (and sometimes even slightly better) with uncorrected reads, highlighting its robustness to noisy ONT reads.
Based on these observations, we select Flye as the assembler to generate the gold standard assemblies in our evaluations, given its ability to handle noisy ONT reads without the need for error correction and its consistent performance across different datasets. Therefore, we use Flye to construct the gold standard assemblies from the basecalled reads in our study.
Table S1: Coverage levels before and after error correction for each dataset.
Dataset Coverage Before Correction Coverage After Correction
D1 E. coli 445$\times$ 240$\times$
D2 Yeast 32$\times$ 12$\times$
D3 Green algae 5.6$\times$ 3.7$\times$
D4 E. coli (R10.4.1) 1796.9$\times$ 338$\times$
C Configuration
C.1 Parameters
[…]
Table S2: Parameters we use in our evaluation for each tool and dataset in mapping.
Tool D1 E. coli D2 Yeast D3 Green Algae D4 E. coli (R10.4.1)
Rawsamble -x ava -t 32 -x ava -t 32 -x ava -t 32 -x ava –r10 -t 32
Minimap2 -x ava-ont –for-only -t 32
Dorado CPU (Fast) basecaller -x cpu fast
Dorado CPU (HAC) basecaller -x cpu hac
Dorado GPU (HAC) basecaller hac
Table S3: Corresponding parameters of presets (-x) in Rawsamble.
Preset Corresponding parameters Usage
ava-viral -e 6 -q 4 -w 0 –sig-diff 0.45 –fine-range 0.4 –min-score 20 –min-score2 30 –min-anchors 5 Viral genomes
–min-mapq 5 –bw 1000 –max-target-gap 2500 –max-query-gap 2500 –chain-gap-scale 1.2 –chain-skip-scale 0.3
ava -e 8 -q 4 -w 3 –sig-diff 0.45 –fine-range 0.4 –min-score 40 –min-score2 75 Default case
–min-anchors 5 –min-mapq 5 –bw 5000 –max-target-gap 2500 –max-query-gap 2500
C.2 Versions
Supplementary Table S4 shows the version and the link to these corresponding versions of each tool we use in our experiments.
Table S4: Versions of each tool and library.
Tool Version Link to the Source Code
Rawsamble 2.1 https://github.com/CMU-SAFARI/RawHash/releases/tag/v2.1
Minimap2 2.24 https://github.com/lh3/minimap2/releases/tag/v2.24
Dorado 0.9.6 (for R9.4) https://github.com/nanoporetech/dorado/releases/tag/v0.9.6
Dorado 1.0.2 (for R10.4.1) https://github.com/nanoporetech/dorado/releases/tag/v1.0.2
Miniasm 0.3-r179 https://github.com/lh3/miniasm/releases/tag/v0.3
Rawasm main https://github.com/CMU-SAFARI/rawasm
Flye 2.9.5 https://github.com/mikolmogorov/Flye/releases/tag/2.9.5
HERRO 0.1 https://github.com/lbcb-sci/herro
[…]